OpenStudy (anonymous):

The altitude in feet of a projectile t seconds after the launch is given by S(t)=-16t^2+144t+32. Find the maximum height reached by the projectile.

OpenStudy (kropot72):

First differentiate S with respect to t. Can you do this step?

OpenStudy (anonymous):

s speed t is time

OpenStudy (kropot72):

In this question S is defined as altitude in feet

OpenStudy (anonymous):

So we find S' by a basic application of the power rule, resulting in S'(t)=-32t+144. We then find a critical point of the function by setting S'(t)=0, so 32t=144, and t=4.5. We then find the maximum height, or critical value, by substituting this value for t back into S(t), which I will leave to you.

OpenStudy (anonymous):

why is it -32t not -16

OpenStudy (anonymous):

The derivative of \(ax^n\) is \(nax^{n-1}\). So in this case, \(2\cdot-16t\).

OpenStudy (anonymous):

gotcha

OpenStudy (campbell_st):

its easy... find the line of symmetry t = -b/2a t = -144/(2 x 16) max height when t = 4.5 seconds alternatively differentiate and solve for t this is the stationary point -32t + 144 = 0 t = 4.5 same result

OpenStudy (campbell_st):

oops its -144/(2 x -16) gives t = 4.5

OpenStudy (anonymous):

s=79.11

OpenStudy (anonymous):

thats how high it will reach

OpenStudy (campbell_st):

I got 356 feet

OpenStudy (anonymous):

\[ -16(4.5^2)+144(4.5)+32=356 \]

OpenStudy (anonymous):

so dont put 4.5 in with the s in front s(t) thats how i came up with 79.11 divided 356 by 4.5

OpenStudy (anonymous):

so the right answer is 356 feet

OpenStudy (anonymous):

The correct answer is 356.

OpenStudy (anonymous):

thank you cambell st nbouscal and kropot72